OpenStudy (anonymous):

square root of 24x^6y^-4

OpenStudy (anonymous):

\[\sqrt{24x^6y^4}=2x^3y^2\sqrt{6}\]

OpenStudy (anonymous):

(2x^3y^-2) sqrt(6)